# Ecosystem-level Analysis of Deployed Machine Learning Reveals Homogeneous Outcomes

### v1

@@ A.1 Annotator disagreement
Figure 7: We stratify ecosystem-level analysis on fer+ by instance-level annotator disagreement – which we take as a proxy for the ambiguity inherent to the input instance – and plot the difference between observed and expected rates for each instance subset. We observe profile polarization for all subsets of data, regardless of the level of annotator disagreement on the instance.
@@
The test set of fer+ provided in HAPI contains instances with disagreement percentages ranging from 0% to 50%. We stratify on the disagreement percentage of the instances and compare expected and observed ecosystem-level outcomes for each subset of instances.
Polarized profiles manifests, regardless of annotator disagreement.
In Figure 7, we find that polarized profiles appear for all levels of annotator disagreement. While more ambiguous examples exhibit higher model error rates and systemic failure rates, the observed rate of polar outcomes exceeds the expected rate of polar outcomes for all instance subsets. This suggests that instance-level ambiguity does not (fully) explain the existence of polarized profiles — at least in fer+.
Figure 8: Interpretation of the relationship between instance-level annotator disagreement and profile polarization depends on if the strength of the effect is quantified as the ratio between observed and expected rates (left plot) or as the difference between observed and expected rates (right plot).
@@
In light of the observed existence of profile polarization across all levels of disagreement, we now examine how the intensity of this effect varies with annotator disagreement. The relationship between annotator disagreement and the intensity of profile polarization depends on the quantification method used to measure profile polarization intensity. When quantifying the intensity as the difference between observed and expected rates, the effect becomes more pronounced as disagreement increases. However, when considering the intensity as the ratio between observed and expected rates, as in the homogenization metric introduced by Bommasani et al. [2022], the effect is most pronounced at lower levels of disagreement.
@@ A.2 Human accuracy
Figure 9: Profile polarization exists for the subset of instances where both dermatologists misclassify the image but doesn’t exist for the subset where both dermatologists correctly classify the image.
Each instance of the DDI dataset — a medical imaging skin lesion dataset — contains predictions from 2 models and 2 humans with the ground truth generated from an external process (in this case, a biopsy of the lesion). To understand how human-perceived difficulty relates to profile polarization in models, we stratify instances on the dermatologist accuracy of that instance (each instance has a dermatologist accuracy of 0%, 50%, or 100%) and examine profile polarization for each subset of instances.
Model outcomes are polarized for instances that dermatologists fail at and unpolarized for those that dermatologists succeed at.
In Figure 9, we find that model outcomes exhibit profile polarization for the subset of instances that both human dermatologists fail at to an even greater extent than observed at the population level. In contrast, the observed and expected distributions match each other (meaning there is no profile polarization, anti-polarization or other form of deviation between the distributions) for the subset of instances that both dermatologists get correct. Note that in the instances that both dermatologists fail at, both models systemically fail more than expected, but both also succeed more than expected. While the correlation between human systemic failures and model systemic failures seems intuitive, it’s less clear why models jointly succeed more than expected on human systemic failures.
Given these surprising and intuitive findings, we encourage future work to explain what features of these instances lead models to pattern together. We speculate that these instances lack sufficient informational cues, prompting the models to place excess emphasis on a narrow selection of features, consistent with the literature of machine learning models being susceptible to spurious correlations [e.g. Sagawa et al., 2020]. At face value, these results would suggest that human-level difficulty can be predictive of model outcome polarization, but we emphasize two caveats: (i) the DDI dataset’s sample size is limited (155 instances of dermatologists both failing and 356 instances of dermatologists both succeeding) and (ii) we rely on the judgments of just two human domain experts, meaning the findings may not generalize to larger annotator pools of non-experts (as are common in many NLP or computer vision datasets). In general, we caution against overgeneralizing this finding. We provide it as an initial foray into understanding example difficulty in a unique setting where we have structure to proxy human-perceived difficulty.
@@ A.3 More expressive theoretical models
Finally, while we consistently find that observed ecosystem-level behaviors yield more homogeneous outcomes and polarized profiles than the instance-level independent model predictions would predict, we might intuit there exists instance-level structure that should be encoded into the prior. Therefore, we consider a theoretical framework to encode richer priors on what we might expect of models. As a simple model, we will assume some instances are universally ‘hard’, meaning all models will perform worse on average across these instances, and others are conversely ‘easy‘, meaning all models will perform better on average across these instances. [^7] As before, we note that there is not a universal standard according to which examples can be considered ‘hard.’ Some examples are easy for humans but hard for models; other examples are easy for models but challenging for humans; and still other examples are challenging for some models but not others. The three hypotheses we consider in this section explore these observer-relative dimensions of hardness.
We use two parameters ($\alpha,\Delta$) to parameterize this model, thereby adjusting the ‘expected’ rate that we calculate for ecosystem-level outcomes. $\alpha$ specifies the composition of ‘hard’ vs ‘easy’ instances in a dataset and $\Delta$ controls how much harder or easier the hard or easy examples, respectively, are expected to be. Concretely, $\alpha$ fraction of instances are categorized as difficult and the remaining $1-\alpha$ are easy. A model’s failure rate $\bar{f_{j}}$ over all examples scales to $\bar{f_{j}}^{\text{hard}}=(1+\Delta)\bar{f_{j}}$ on hard examples and $\bar{f_{j}}^{\text{easy}}=\left(1-\frac{\alpha\Delta}{1-\alpha}\right)\bar{f_{j}}$ on easy examples.
The distribution of the expected number of model failures $t\in\{0,\dots,k\}$ follows a weighted sum of two Poisson-Binomial distributions parameterized by the scaled hard $\bar{f_{j}}^{\text{hard}}$ and easy $\bar{f_{j}}^{\text{easy}}$ model error rates.
$P_{\text{expected}}^{\text{hard}}(t~{}\text{failures})$ $=(\alpha)\text{Poisson-Binomial}(\bar{f_{1}}^{\text{hard}},\dots,\bar{f_{k}}^{\text{hard}})[t]$ (3)
$P_{\text{expected}}^{\text{easy}}(t~{}\text{failures})$ $=(1-\alpha)\text{Poisson-Binomial}(\bar{f_{1}}^{\text{easy}},\dots,\bar{f_{k}}^{\text{easy}})[t]$ (4)
$P_{\text{expected}}(t~{}\text{failures})$ $=P_{\text{expected}}^{\text{hard}}[t]+P_{\text{expected}}^{\text{easy}}[t]$ (5)
Figure 10: Observed and expected distribution of ecosystem-level outcomes for the $(\alpha,\Delta)$ combination that yields the lowest L1 distance for each dataset.
@@
We utilize this framework to identify which $(\alpha,\Delta)$ combinations generate expected distributions that recover the observed distributions in the HAPI datasets. We perform a grid search for $\alpha\in[0.1,0.5]$ with a step size of $0.1$ and $\Delta\in[0.2,5]$ with a step size of $0.2$. Note that certain $(\alpha,\Delta)$ combinations can result in invalid error rates depending on the original error rates of models – i.e. when $\left(1-\frac{\alpha\Delta}{1-\alpha}\right)\bar{f_{j}}<0$.
Figure 11: L1 distance between observed and expected distributions for all datasets as a function of $\alpha$ and $\Delta$.
For each dataset, we identify the $(\alpha,\Delta)$ combination that minimizes the L1 distance (equivalently the total variation distance) between the observed and expected distributions. In Figure 10, we visualize the observed and expected distributions for the distance-minimizing $(\alpha,\Delta)$ pair for each dataset, and in Figure 11 we plot the L1 distance for each dataset as a function of $\alpha$ and $\Delta$. The majority of the best $\alpha$ values are $0.2$ or $0.3$ while the $\Delta$ values can range from $1$ to $4$ – indicating that the $\bar{f_{j}}^{\text{hard}}$ can be up to 5x higher than $\bar{f_{j}}$.
@@ B.1 Data
B.2 Leader Following Effects in Profile Polarization
@@
One consequence of profile polarization is that it concentrates failures on the same users, so a user who interacts with a model and experiences a failure from that model is now more likely to experience a failure from every other model in the ecosystem. To quantify the strength of this effect, we examine how the probability of a user experiencing each outcome profile changes after that user experiences one failure from a model.
In Figure 12, we find that, consistent with the observed profile polarization in HAPI, observing a single model failure significantly increases the probability that the user will now experience failures from all alternative models in the ecosystem; however, we also find that the strength of this effect is strongly graded by the accuracy of the model that we initially observe a failure for. Upon observing the most accurate model in the ecosystem fail for a user, the probability of that user experiencing a systemic failure increases by 37% – whereas it only increases by around 16% when we observe the least accurate model fail.
This result suggests that instances that the most accurate model fails on are likely to be failed by the less accurate models as well. This ‘leader following’ phenomenon has implications for users in a model ecosystem: users failed by the most accurate model likely have drastically fewer options for alternative models that could work for them.
@@ B.3 Model improvement analysis is insensitive to threshold
In § 4, we study how the improvement of a single model, in the sense that it becomes more impact, manifests at the ecosystem-level. To define improvement, we set the (slightly arbitrary) threshold that the model’s accuracy improve by $0.5\%$, which we found to be large enough to be material to a model’s performance while small enough to capture most model improvements. In Figure 13, we plot the outcome profile distribution in the observed improvements set against the distribution over the potential improvements (as in 3(b)) for 6 different thresholds of change. Across all thresholds, the patterns we discuss in § 4 hold: namely, models consistently under-improve on systemic failures. This confirms that the findings and qualitative understanding we present is not particularly sensitive to the exact value of this threshold.
@@ C.2 Analyses are insensitive to including/excluding HAM10k
((a)) Model outcomes when including HAM10K yield even more polarized profiles on DDI than in Figure 5
@@
Figure 15: We replicate the graphs in Figure 5 and Figure 6 but with the inclusion of HAM10K. Profile polarization and racial disparities in models are even more pronounced when including HAM10k.
@@
We decided to remove HAM10k because the pattern of near-universal negative predictions does not reflect model behavior we would expect of models deployed in clinical settings. Namely, even if deployed, the structure of the model errors are not particularly interesting and are largely predictable (in direction). Beyond these fundamental reasons for excluding the model, we also removed the model for reasons unique to our analysis. Including the model would have introduced an explicit class correlation in systemic failures (i.e. almost all systemic failures would be malignant instances and none would be benign instances) and would have complicated the comparison with humans (since there would be three models vs. two humans).
To confirm that our findings hold independent of this choice, in Figure 15, we replicate Figure 5 and Figure 6 but with outcomes from HAM10k included. We find that inclusion of HAM10k exacerbates the profile polarization of model outcomes and the racial disparities in model outcomes.

### v2

@@ A.1 Annotator disagreement
Figure 7: We stratify ecosystem-level analysis on fer+ by instance-level annotator disagreement – which we take as a proxy for the ambiguity inherent to the input instance – and plot the difference between observed and baseline rates for each instance subset. We observe homogeneous outcomes for all subsets of data, regardless of the level of annotator disagreement on the instance.
@@
The test set of fer+ provided in HAPI contains instances with disagreement percentages ranging from 0% to 50%. We stratify on the disagreement percentage of the instances and compare baseline and observed ecosystem-level outcomes for each subset of instances.
Homogenous Outcomes manifest regardless of annotator disagreement.
In Figure 7, we find that homogenous outcomes appear for all levels of annotator disagreement. While more ambiguous examples exhibit higher model error rates and systemic failure rates, the observed rate of homogeneous outcomes exceeds the baseline rate of homogeneous outcomes for all instance subsets. This suggests that instance-level ambiguity does not (fully) explain the existence of homogenous outcomes— at least in fer+.
Figure 8: Interpretation of the relationship between instance-level annotator disagreement and homogeneous outcomes depends on if the strength of the effect is quantified as the ratio between observed and baseline rates (left plot) or as the difference between observed and baseline rates (right plot).
@@
In light of the observed existence of homogeneous outcomes across all levels of disagreement, we now examine how the intensity of this effect varies with annotator disagreement. The relationship between annotator disagreement and the intensity of homogeneous outcomes depends on the quantification method used to measure homogeneous outcomes intensity. When quantifying the intensity as the difference between observed and baseline rates, the effect becomes more pronounced as disagreement increases. However, when considering the intensity as the ratio between observed and baseline rates, as in the homogenization metric introduced by Bommasani et al. [2022], the effect is most pronounced at lower levels of disagreement.
@@ A.2 Human accuracy
Figure 9: Homogeneous outcomes exists for the subset of instances where both dermatologists misclassify the image but doesn’t exist for the subset where both dermatologists correctly classify the image.
Each instance of the DDI dataset — a medical imaging skin lesion dataset — contains predictions from 2 models and 2 humans with the ground truth generated from an external process (in this case, a biopsy of the lesion). To understand how human-perceived difficulty relates to homogeneous outcomes, we stratify instances on the dermatologist accuracy of that instance (each instance has a dermatologist accuracy of 0%, 50%, or 100%) and examine homogeneous outcomes for each subset of instances.
Model outcomes are homogenous for instances on which dermatologists fail and heterogeneous for on which dermatologists succeed.
In Figure 9, we find that model outcomes exhibit homogeneous outcomes for the subset of instances that both human dermatologists fail at to an even greater extent than observed at the population level. In contrast, the observed and baseline distributions match each other (meaning there is no homogeneous outcomes, heterogeneity or other form of deviation between the distributions) for the subset of instances that both dermatologists get right. Note that in the instances that both dermatologists fail at, both models systemically fail more than the baseline, but both also succeed more than the baseline. While the correlation between human systemic failures and model systemic failures seems intuitive, it’s less clear why models jointly succeed more than the baseline on human systemic failures.
Given these surprising and unintuitive findings, we encourage future work to explain what features of these instances lead models to pattern together. We speculate that these instances lack sufficient informational cues, prompting the models to place excess emphasis on a narrow selection of features, consistent with the literature of machine learning models being susceptible to spurious correlations [e.g. Sagawa et al., 2020]. At face value, these results would suggest that human-level difficulty can be predictive of outcome homogeneity, but we emphasize two caveats: (i) the DDI dataset’s sample size is limited (155 instances of dermatologists both failing and 356 instances of dermatologists both succeeding) and (ii) we rely on the judgments of just two human domain experts, meaning the findings may not generalize to larger annotator pools of non-experts (as are common in many NLP or computer vision datasets). In general, we caution against overgeneralizing this finding. We provide it as an initial foray into understanding example difficulty in a unique setting where we have data that supports an approximation of human-perceived difficulty.
@@ A.3 More expressive theoretical models
Finally, while we consistently find that observed ecosystem-level behaviors yield more homogeneous outcomes than the instance-level independent model predictions would predict, we might intuit there exists instance-level structure that should be encoded into the prior. Therefore, we consider a theoretical framework to encode richer priors on what we might expect of models. As a simple model, we will assume some instances are universally ‘hard’, meaning all models will perform worse on average across these instances, and others are conversely ‘easy‘, meaning all models will perform better on average across these instances.
As before, we note that there is not a universal standard according to which examples can be considered ‘hard.’ Some examples are easy for humans but hard for models; other examples are easy for models but challenging for humans; and still other examples are challenging for some models but not others. The three hypotheses we consider in this section explore these observer-relative dimensions of hardness.
We use two parameters ($\alpha,\Delta$) to parameterize this model, thereby adjusting the baseline rate that we calculate for ecosystem-level outcomes. $\alpha$ specifies the composition of ‘hard’ vs ‘easy’ instances in a dataset and $\Delta$ controls how much harder or easier the hard or easy examples, respectively, are expected to be. Concretely, $\alpha$ fraction of instances are categorized as difficult and the remaining $1-\alpha$ are easy. A model’s failure rate $\bar{f_{j}}$ over all examples scales to $\bar{f_{j}}^{\text{hard}}=(1+\Delta)\bar{f_{j}}$ on hard examples and $\bar{f_{j}}^{\text{easy}}=\left(1-\frac{\alpha\Delta}{1-\alpha}\right)\bar{f_{j}}$ on easy examples.
The distribution of the baseline number of model failures $t\in\{0,\dots,k\}$ follows a weighted sum of two Poisson-Binomial distributions parameterized by the scaled hard $\bar{f_{j}}^{\text{hard}}$ and easy $\bar{f_{j}}^{\text{easy}}$ model error rates.
$P_{\text{baseline}}^{\text{hard}}(t~{}\text{failures})$ $=(\alpha)\text{Poisson-Binomial}(\bar{f_{1}}^{\text{hard}},\dots,\bar{f_{k}}^{\text{hard}})[t]$ (3)
$P_{\text{baseline}}^{\text{easy}}(t~{}\text{failures})$ $=(1-\alpha)\text{Poisson-Binomial}(\bar{f_{1}}^{\text{easy}},\dots,\bar{f_{k}}^{\text{easy}})[t]$ (4)
$P_{\text{baseline}}(t~{}\text{failures})$ $=P_{\text{baseline}}^{\text{hard}}[t]+P_{\text{baseline}}^{\text{easy}}[t]$ (5)
Figure 10: Observed and baseline distributions of ecosystem-level outcomes for the $(\alpha,\Delta)$ combination that yields the lowest L1 distance for each dataset.
@@
We utilize this framework to identify which $(\alpha,\Delta)$ combinations generate baseline distributions that recover the observed distributions in the HAPI datasets. We perform a grid search for $\alpha\in[0.1,0.5]$ with a step size of $0.1$ and $\Delta\in[0.2,5]$ with a step size of $0.2$. Note that certain $(\alpha,\Delta)$ combinations can result in invalid error rates depending on the original error rates of models – i.e. when $\left(1-\frac{\alpha\Delta}{1-\alpha}\right)\bar{f_{j}}<0$.
Figure 11: L1 distance between observed and baseline distributions for all datasets as a function of $\alpha$ and $\Delta$.
For each dataset, we identify the $(\alpha,\Delta)$ combination that minimizes the L1 distance (equivalently the total variation distance) between the observed and baseline distributions. In Figure 10, we visualize the observed and baseline distributions for the distance-minimizing $(\alpha,\Delta)$ pair for each dataset, and in Figure 11 we plot the L1 distance for each dataset as a function of $\alpha$ and $\Delta$. The majority of the best $\alpha$ values are $0.2$ or $0.3$ while the $\Delta$ values can range from $1$ to $4$ – indicating that the $\bar{f_{j}}^{\text{hard}}$ can be up to 5x higher than $\bar{f_{j}}$.
@@ B.1 Data
B.2 Leader Following Effects in Systemic Failure
@@
One consequence of homogeneous outcomes is that it concentrates failures on the same users, so a user who interacts with a model and experiences a failure from that model is now more likely to experience a failure from every other model in the ecosystem. To quantify the strength of this effect, we examine how the probability of a user experiencing each outcome profile changes after that user experiences one failure from a model.
In Figure 12, we find that, consistent with the observed homogeneous outcomes in HAPI, observing a single model failure significantly increases the probability that the user will now experience failures from all other models in the ecosystem. However, we also find that the strength of this effect is strongly graded by the accuracy of the model for which we initially observe a failure. Upon observing the most accurate model in the ecosystem fail for a user, the probability of that user experiencing a systemic failure increases by 37% – whereas it only increases by around 16% when we observe the least accurate model fail.
This result suggests that instances that the most accurate model fails on are likely to be failed by the less accurate models as well. This ‘leader following’ phenomenon has implications for users in a model ecosystem: users failed by the most accurate model likely have few options for alternative models that could work for them.
@@ B.3 Model improvement analysis is insensitive to threshold
In § 4, we study how the improvement of a single model, in the sense that it becomes more impact, manifests at the ecosystem-level. To define improvement, we set the (slightly arbitrary) threshold that the model’s accuracy improve by $0.5\%$, which we found to be large enough to be material to a model’s performance while small enough to capture most model improvements. In Figure 13, we plot the outcome profile distribution in the observed improvements set against the distribution over the potential improvements (as in 4(b)) for 6 different thresholds of change. Across all thresholds, the patterns we discuss in § 4 hold: namely, models consistently under-improve on systemic failures. This confirms that the findings and qualitative understanding we present is not particularly sensitive to the exact value of this threshold.
@@ B.4 Net Improvements
B.5 When Models Get Worse
As a related question to what we examine in 4(b), we examine what happens to ecosystem-level outcomes when a model gets worse. We find that, when models get worse, they disproportionately introduce new systemic failures into the system by ’over-declining’ on instances that other models were already failing for. This further highlights how single-model measurements often fails to align with ecosystem-level outcomes.
Figure 15: The distribution of outcome profiles for all year-over- year model declines across all datasets
@@ C.2 Analyses are insensitive to including/excluding HAM10k
((a)) Model outcomes when including HAM10K yield even more homogenous outcomes on DDI than in Figure 5
@@
Figure 16: We replicate the graphs in Figure 5 and Figure 6 but with the inclusion of HAM10K. Homogeneous outcomes and racial disparities in models are even more pronounced when including HAM10k.
@@
We decided to remove HAM10k because the pattern of near-universal negative predictions does not reflect model behavior we would expect of models deployed in clinical settings. Namely, even if deployed, the structure of the model errors are not particularly interesting and are largely predictable (in direction). Beyond these fundamental reasons for excluding the model, we also removed the model for reasons unique to our analysis. Including the model would have introduced an explicit class correlation in systemic failures (i.e. almost all systemic failures would be malignant instances and none would be benign instances) and would have complicated the comparison with humans since there would be three models but only two humans.
To confirm that our findings hold independent of this choice, in Figure 16, we replicate Figure 5 and Figure 6 but with outcomes from HAM10k included. We find that inclusion of HAM10k exacerbates the homogeneous outcomes of model outcomes and the racial disparities in model outcomes.